OpenStudy (azureflame241):

How do i simplify log(1/100)

terenzreignz (terenzreignz):

I'm quite certain 1/100 is ten raised to something...

OpenStudy (anonymous):

-2 is the answer use calculator :D

OpenStudy (azureflame241):

thanks jacal, but i guess i should of specified my question as "to be simplified w/o cal"

OpenStudy (anonymous):

dont call me jacal call me aila please

OpenStudy (anonymous):

its hard to find wthout using calculator :D

OpenStudy (phi):

log without the base specified generally means \( \log_{10} \) read log base 10 It helps to know \[ \log(10^a) = a\] when you see log, think of two things: the base, and the exponent of the base

OpenStudy (phi):

for your problem, it would be convenient to write 1/100 as 10 to some power hopefully you know 10*10= 100, or, using exponents, 10^2 = 100 so 1/100 is the same as 1/10^2 that is close to what you want. but it is "upside down" However, you should remember that \[ \frac{1}{10^a} = 10^{-a} \] that is a good rule to remember. use that rule to write 1/10^2 as 10^-2 now what is \[ \log(10^{-2} )\]

OpenStudy (azureflame241):

its -2Log10

OpenStudy (phi):

and what is \[ \log(10^1) \]

OpenStudy (azureflame241):

1?

OpenStudy (anonymous):

correct

OpenStudy (phi):

the simple rule is \[ \log(10^a) = a \] so log(10^1) is 1 and -2*1 = -2 notice you could use the rule to get log(10^-2) = -2 immediately or you could do what you did

OpenStudy (azureflame241):

oh ok Thanks!

OpenStudy (anonymous):

click the best response please